Sergio is 70 years younger than Elizabeth. 8 years ago, Elizabeth’s age was 3 times Sergio’s age. How old is Sergio now?
Answer by VFBundy(438)

Sergio's age = x
Elizabeth's age = x + 70

"8 years ago, Elizabeth’s age was 3 times Sergio’s age."

(x + 70) - 8 = 3(x - 8)

x + 62 = 3x - 24

86 = 2x

x = 43

Sergio's age = x = 43
Elizabeth's age = x + 70 = x + 43 = 113

==========================================

Check it:

"Sergio is 70 years younger than Elizabeth."

Sergio = 43; Elizabeth = 113. 113 minus 43 = 70. [CHECKS]

"8 years ago, Elizabeth’s age was 3 times Sergio’s age."

Sergio = 35; Elizabeth = 105. 35 times 3 = 105. [CHECKS]